# SurgRIPE challenge: Benchmark of Surgical Robot Instrument Pose Estimation

### v1

@@ 4 Challenge submissions
For object detection, YOLOv5 Ultralytics (2023) was used, which provides an effective solution for detecting 2D bounding boxes of object instances. Given the small number of training images, the training is based on a model pre-trained on the COCO dataset Lin et al. (2014). More specifically, their medium-size model trained on higher resolution images, YOLOv5m6, was selected. For training, we extract the bounding boxes of the ground truth masks to provide labels. For validation, we randomly sample 20% from the training set. Empirically decided, we train this model for 300 epochs with a batch size of 8. For data augmentation and other hyper-parameters, we follow the default, recommended settings of the original implementation of YOLOv5.
For 6D object pose estimation, we employ SurfEmb Haugaard and Buch (2022). Unlike direct pose regression methods that require a high number of training samples capturing a large variation of poses of the object, SurfEmb uses surface coordinates of the object as the training target. This enables an efficient training scheme for the relatively low number of provided training samples. Furthermore, SurfEmb employs metric learning across a learned implicit representation of the object surface and their projections on the 2D images enabling an effective capture of the unknown and complex symmetries of the target objects. For training SurfEmb, the default settings proposed by the authors were followed. One single ResNet18 He et al. (2016) encoder was employed, for both surgical instruments (LND and MBF), coupled with separate decoders for each instrument. For the implicit representations of the object surfaces, we use a SIREN-based Sitzmann et al. (2020) multiple-layer perceptron (MLP) for each instrument. Given the low number of training samples, we separate only 20 samples for each instrument from the training set for validation. We train our model for 10 epochs with a batch size of 16 and employ the intensity and geometric data augmentations utilized in the original implementation. We use the default values of the other hyper-parameters.
@@
During inference, first YOLOv5 Ultralytics (2023) detects an instance of the targeted surgical tool on the input image of size 960 x 540. This predicted bounding box is then used to crop a square patch around its center and resize it to 224 x 224. The convolutional neural network (CNN) part of our SurfEmb model takes in the cropped image along with the predicted class label and estimates 2D dense descriptors and the binary object mask. In addition, its MLP component predicts 3D descriptors on the densely sampled surface points of the targeted object 3D model. The 2D-3D matches are recovered through cosine similarity between the descriptors of the two domains. Since the surgical instruments have certain symmetries that can cause spatial ambiguities, we follow SurfEmb’s proposed multiple-hypothesis-based pose estimation strategy that creates numerous 2D-3D correspondence subsets and employs them within AP3P Ke and Roumeliotis (2017) and select the maximum scoring pose considering the object segmentation mask, correspondence distribution and the probability of the visible surface coordinates. Finally, the predicted pose is refined through a render and compare strategy using BFGS optimizing for the maximization of the 2D-3D correspondence score.
@@
This method improves PVNet Peng et al. (2019) by retaining its core pose estimation mechanism while introducing targeted data augmentation strategies and a de-glare algorithm to improve the algorithm’s handling of occlusions and reflective surfaces in surgical environments.
@@ 4.3.2 Segmentation and Regression models
DeepLabV3+ Chen et al. (2018) is used for the segmentation step. We adopt ConvNeXT Liu et al. (2022) as the regression backbone. With the feature produced by the backbone, a pose vector and a depth map are predicted. This set of multi-tasking can enhance the pose estimation. The ground truth depth for the instrument region can be generated using the ground truth pose and the instrument model. The upper head is adopted as the head for depth map generation. The depth map is not used for refinement in post-processing step.
@@ 4.3.3 Data pre-process
As mentioned above, a copy-paste method is adopted. We use the segmentation mask to extract only instrument pixels and superimpose them onto images of different scenes. A background picture and a foreground picture are paired only if their bounding boxes do not overlap
@@
The EfficientPose Bukschat and Vetter (2020) architecture extends the capabilities of the EfficientNet model by incorporating two additional sub-networks for precise pose estimation. EfficientPose utilizes the EfficientNet as the backbone network and includes the Bi-directional Feature Pyramid Network (BiFPN) structure. This extension is tailored to perform efficient and accurate detection, achieving pose estimation performance using RGB images only.
@@
This framework’s network architecture follows an encoder-decoder structure, a design commonly employed for pose estimation tasks that consistently yields favourable outcomes. In the encoding phase, we adopt a modified version of ResNet-18 and incorporate a transformer encoder proposed by Shaker et al. Shaker et al. (2023). The Swift-Former Shaker et al. (2023), bolstered by its efficient additive attention mechanism, empowers the model to discern contextual relationships across distant image regions. It can enhance our model’s ability to extract global context information by effectively integrating contextual cues from different regions. We believe this characteristic is valuable in this surgical tool pose estimation scenario, given the elongated and intricate nature of surgical tools.
Similar to PVNet Peng et al. (2019), we train the network in a supervised manner to learn a semantic mask and a vertex map, which is a pixel-wise representation of the keypoints. The vertex map is stored as two channels: one for the $dx$ values and another for the $dy$ values across the entire image. Each keypoint has its corresponding vertex map with $dx$ and $dy$ values, resulting in a total of $2\times k$ channels in the vertex branch of the network, with dimensions $H\times W$. Using the estimated semantic mask and unit vectors as inputs, we employ a RANSAC-based voting approach to generate potential keypoint hypotheses. After obtaining the estimated 2D keypoints that correspond to the sampled 3D keypoints, along with the camera intrinsics from the dataset, we use the solvePnP function in OpenCV to solve for the pose.
@@ 4.6.1 PVNet
PVNet Peng et al. (2019) is a state-of-the-art object pose estimation method which utilises pixel-wise unit vectors to estimate keypoints for keypoint detection. Then the detected keypoints can be used to solve a Perspective-n-Point (PnP) problem for pose estimation. A Pixel-wise Voting Network is introduced to localise pixel-wise unit vectors, then uses these vectors for keypoint voting using differentiable RANSAC method. This unit-vector representation is flexible for localising occlusion and truncated parts of objects, which is common in surgical scenes.
@@ 5.1.1 Methods
Figure 8: Visualisation of Multi Losses in UOL Method
Loss Function: We used a multi-task loss function that included points, projection, translation and contrastive loss terms. For the translation loss, the RMSE was calculated between predicted and ground truth translation parameters and minimised. For the points loss, the 3D model of the tool joint was transformed by both the predicted and ground truth pose, and then the average L1 distance between corresponding points was calculated. This loss enforced consistency in 3D space and is shown in Fig. 8(b). The projection loss term enforces consistency between the 3D and 2D spaces, as seen in Fig. 8(a). The 3D tool is transformed according to the ground truth and predicted poses, then projected into the 2D image plane using the camera intrinsics. The concave hull of the 2D points was then used to obtain a projected binary segmentation mask of the tool head and the dice loss was minimised between this and the ground truth segmentation mask.
@@ 6 Challenge results
The challenge results for each subset are presented in Tables 3- 6. We noticed that if the pose estimation fails in one frame, it will cause an outlier with a large translation and rotation error. This makes the average translation and rotation error of several methods heavily affected by these extreme error values. Therefore, to rank the competing methods, we mainly focus on the ADD, Accuracy–ADD threshold Curve and the Avg Acc as defined in Section 3.2.
Team ADD (10% diameter) ↑ Avg Acc (0-5 MM)↑ Translation Error (mm) ↓ Rotation Error (degree) ↓
@@
UOL 0.0617 0.2354 4.3196 18.5886
PVNet 0.1930 0.2866 46.7894 52.4488
Table 3: Evaluation on LND Test Without Occlusion subset.
Team ADD (10% diameter) ↑ Avg Acc (0-5 MM)↑ Translation Error (mm) ↓ Rotation Error (degree) ↓
@@
UOL 0.0191 0.0572 7.9181 10.7929
PVNet 0.3589 0.4198 3.5265 25.8501
Table 4: Evaluation on MBF Test Without Occlusion subset.
@@ 6.1 Instrument Pose Estimation Without Occlusion
The instrument pose estimation performance without occlusion was evaluated across two test subsets namely, LND Test Without Occlusion and MBF Test Without Occlusion using the ADD (10% diameter), Avg Acc (0-5 MM), Translation Error, and Rotation Error as shown in Table 3 and Table 4. Overall, IGTUM outperformed the other methods, consistently achieving the highest accuracy (ADD and Avg Acc) and the lowest errors (Translation and Rotation) across both subsets, indicating its robustness and effectiveness. It achieved 36.06% ADD on average and 45.48% Avg Acc for both subsets. MVL_3S demonstrated good performance in the MBF Test Without Occlusion set (Table 4), with competitive ADD and Avg Acc scores and overall lower translation and rotation errors. It achieved 17.19% ADD on average and 26.38% Avg Acc for both subsets. TUDU exhibited moderate accuracy overall with 10.09% ADD on average and 20.15% Avg Acc for both subsets. UOL achieved 4.04% ADD on average and 14.63% Avg Acc for both subsets. To be noticed, MVL_3S and TUDU achieved low translation errors but higher rotation errors. ICL, EUT and PVNet demonstrated strong performance in terms of ADD and Avg Acc scores but higher errors in translation and rotation error. EUT achieved 17.94% ADD on average and 24.06% Avg Acc while ICL obtained 22.92% ADD on average and 29.52% Avg Acc for both subsets. Fig. 9(a) and Fig. 9(b) illustrate the Accuracy curve with respect to different ADD thresholds for the examined subsets. It can be noticed that the accuracy of IGTUM has a significant gap from the other compared methods for the subset LND Test Without Occlusion while IGTUM, ICL, MVL_3S and PVNet have comparable accuracy for the MBF Test Without Occlusion subset.
@@
Figure 9: Accuracy–ADD threshold Curves for test data in Testsets Without Occlusion
@@ 6.2 Instrument Tracking with Occlusion
Considering that pose estimation with occlusion is quite challenging, the performance of all methods under occlusion shows a significant drop as shown in Tables. 5 and Tables. 6. Across the LND Test With Occlusion and MBF Test With Occlusion subsets, IGTUM consistently outperformed the other methods with the highest ADD and Avg Acc scores and the lowest translation and rotation errors. This confirms its robustness and reliability under challenging conditions, such as occlusions. ICL, EUT and PVNet also showed promising performance in terms of ADD and Avg Acc scores, but with higher translation and rotation errors than IGTUM, indicating their limited effectiveness in challenging scenarios. TUDU and MVL_3S had lower ADD and Avg Acc scores across both subsets, coupled with moderate to high translation and rotation error rates, suggesting that these methods are less reliable under occlusion. The above performance trends are verified by the Accuracy-ADD threshold curves shown in Fig. 10(a) and Fig. 10(b).
The performance of the compared methods has been evaluation on the secondary validation metrics, namely proj2d in Table 7 and mmd5 in Table 8. IGTUM still outperformed the other methods with the best proj2d and mmd5 scores. ICL, EUT and PVNet demonstrated the second-best performance in the group since they followed a similar pipeline. UOL, TUDU and MVL_3S had lower proj2d and mmd5 scores across both subsets.
@@
Figure 10: Accuracy–ADD threshold Curves for test data in Testsets With Occlusion
Team ADD (10% diameter) ↑ Avg Acc (0-5 MM)↑ Translation Error (mm) ↓ Rotation Error (degree) ↓
@@
UOL 0.0084 0.0725 8.6025 24.8517
PVNet 0.2731 0.3901 28.0907 17.5543
Table 5: Evaluation on LND Test With Occlusion.
Team ADD (10% diameter) ↑ Avg Acc (0-5 MM)↑ Translation Error (mm) ↓ Rotation Error (degree) ↓
@@
UOL 0.0155 0.0451 14.7409 22.0818
PVNet 0.1731 0.1826 44.6732 31.0920
Table 6: Evaluation on MBF Test With Occlusion.
Team LND Test LND Test MBF Test MBF Test
w/o Occlusion w Occlusion w/o Occlusion w Occlusion
@@
UOL 0.0187 0.0 0.0047 0.0077
PVNet 0.5308 0.8182 0.3487 0.5607
Table 7: Evaluation on proj2d
Team LND Test LND Test MBF Test MBF Test
w/o Occlusion w Occlusion w/o Occlusion w Occlusion
@@
UOL 0.0643 0.0042 0.0669 0.0129
PVNet 0.1903 0.0504 0.4928 0.1757
Table 8: Evaluation on mmd5

### v2

@@ 5 Challenge submissions
For object detection, You Only Look Once (YOLO) v5 (Ultralytics, 2023) was used, which provides an effective solution for detecting 2D bounding boxes of object instances. Given the small number of training images, the training is based on a model pre-trained on the Common Objects in Context (COCO) dataset (Lin et al., 2014). More specifically, their medium-size model trained on higher resolution images, YOLOv5m6, was selected. For training, we extract the bounding boxes of the ground truth masks to provide labels. For validation, we randomly sample 20% from the training set. Empirically decided, we train this model for 300 epochs with a batch size of 8. For data augmentation and other hyper-parameters, we follow the default, recommended settings of the original implementation of YOLOv5.
For 6D object pose estimation, we employ SurfEmb (Haugaard and Buch, 2022). Unlike direct pose regression methods that require a high number of training samples capturing a large variation of poses of the object, SurfEmb uses surface coordinates of the object as the training target. This enables an efficient training scheme for the relatively low number of provided training samples. Furthermore, SurfEmb employs metric learning across a learned implicit representation of the object surface and their projections on the 2D images enabling an effective capture of the unknown and complex symmetries of the target objects. For training SurfEmb, the default settings proposed by the authors were followed. One single Residual Network (ResNet) 18 (He et al., 2016) encoder was employed, for both surgical instruments (LND and MBF), coupled with separate decoders for each instrument. For the implicit representations of the object surfaces, we use a SIREN-based (Sitzmann et al., 2020) multiple-layer perceptron (MLP) for each instrument. Given the low number of training samples, we separate only 20 samples for each instrument from the training set for validation. We train our model for 10 epochs with a batch size of 16 and employ the intensity and geometric data augmentations utilized in the original implementation. We use the default values of the other hyper-parameters.
@@
During inference, first YOLOv5 (Ultralytics, 2023) detects an instance of the targeted surgical tool on the input image of size 960 x 540. This predicted bounding box is then used to crop a square patch around its center and resize it to 224 x 224. The convolutional neural network (CNN) part of our SurfEmb model takes in the cropped image along with the predicted class label and estimates 2D dense descriptors and the binary object mask. In addition, its MLP component predicts 3D descriptors on the densely sampled surface points of the targeted object 3D model. The 2D-3D matches are recovered through cosine similarity between the descriptors of the two domains. Since the surgical instruments have certain symmetries that can cause spatial ambiguities, we follow SurfEmb’s proposed multiple-hypothesis-based pose estimation strategy that creates numerous 2D-3D correspondence subsets and employs them within AP3P (Ke and Roumeliotis, 2017) and select the maximum scoring pose considering the object segmentation mask, correspondence distribution and the probability of the visible surface coordinates. Finally, the predicted pose is refined through a render and compare strategy using Broyden–Fletcher–Goldfarb–Shanno (BFGS) optimizing for the maximization of the 2D-3D correspondence score.
@@
This method improves Pixel-wise Voting Network (PVNet) (Peng et al., 2019) by retaining its core pose estimation mechanism while introducing targeted data augmentation strategies and a de-glare algorithm.
@@
Figure 7: TUDU architecture diagram. The method includes (1) The segmentation branch which segments and crops the instrument. (2) The pose prediction branch where the backbone network predicts the tool pose based on the cropped regions.
@@ 5.3.2 Segmentation and Regression models
DeepLabV3+ (Chen et al., 2018) is used for the segmentation step. We adopt ConvNeXT (Liu et al., 2022) as the regression backbone. With the feature produced by the backbone, a pose vector and a depth map are predicted. This set of multi-tasking can enhance the pose estimation. The ground truth depth for the instrument region can be generated using the ground truth pose and the instrument model. The upper head is adopted as the head for depth map generation. The depth map is not used for refinement in post-processing step.
@@ 5.3.3 Data pre-process
As mentioned above, a copy-paste method is adopted. We use the segmentation mask to extract only instrument pixels and superimpose them onto images of different scenes. A background picture and a foreground picture are paired only if their bounding boxes do not overlap.
@@
The EfficientPose (Bukschat and Vetter, 2020) architecture extends the capabilities of the EfficientNet model by incorporating two additional sub-networks for precise pose estimation. EfficientPose utilizes the EfficientNet as the backbone network and includes the Bi-directional Feature Pyramid Network (BiFPN) structure. This extension is tailored to perform efficient and accurate detection, achieving pose estimation performance using RGB images only.
@@
This framework’s network architecture follows an encoder-decoder structure, a design commonly employed for pose estimation tasks that consistently yields favourable outcomes. In the encoding phase, we adopt a modified version of ResNet-18 and incorporate a transformer encoder proposed by (Shaker et al., 2023). The Swift-Former, bolstered by its efficient additive attention mechanism, empowers the model to discern contextual relationships across distant image regions. It can enhance our model’s ability to extract global context information by effectively integrating contextual cues from different regions. We believe this characteristic is valuable in this surgical tool pose estimation scenario, given the elongated and intricate nature of surgical tools.
Similar to PVNet (Peng et al., 2019), we train the network in a supervised manner to learn a semantic mask and a vertex map, which is a pixel-wise representation of the keypoints. The vertex map is stored as two channels: one for the $dx$ values and another for the $dy$ values across the entire image. Each keypoint has its corresponding vertex map with $dx$ and $dy$ values, resulting in a total of $2\times k$ channels in the vertex branch of the network, with dimensions $H\times W$. Using the estimated semantic mask and unit vectors as inputs, we employ a RANSAC-based voting approach to generate potential keypoint hypotheses. After obtaining the estimated 2D keypoints that correspond to the sampled 3D keypoints, along with the camera intrinsics from the dataset, we use the solvePnP function in OpenCV to solve for the pose.
@@ 5.6.1 PVNet
PVNet (Peng et al., 2019) is a state-of-the-art object pose estimation method which utilises pixel-wise unit vectors to estimate keypoints for keypoint detection. Then the detected keypoints can be used to solve a Perspective-n-Point (PnP) problem for pose estimation. A Pixel-wise Voting Network is introduced to localise pixel-wise unit vectors, then uses these vectors for keypoint voting using differentiable RANSAC method. This unit-vector representation is flexible for localising occlusion and truncated parts of objects, which is common in surgical scenes.
@@ 6.1.1 Methods
Loss Function: We used a multi-task loss function that included points, projection, translation and contrastive loss terms. For the translation loss, the root mean square error (RMSE) was calculated between predicted and ground truth translation parameters and minimised. For the points loss, the 3D model of the tool joint was transformed by both the predicted and ground truth pose, and then the average L1 distance between corresponding points was calculated. This loss enforced consistency in 3D space and is shown in Fig. 8(b). The projection loss term enforces consistency between the 3D and 2D spaces, as seen in Fig. 8(a). The 3D tool is transformed according to the ground truth and predicted poses, then projected into the 2D image plane using the camera intrinsics. The concave hull of the 2D points was then used to obtain a projected binary segmentation mask of the tool head and the dice loss was minimised between this and the ground truth segmentation mask.
@@
Figure 8: Visualisation of Multi Losses in UOL Method.
@@ 7 Challenge results
There were 44 registered teams by the submission deadline. Five teams provided valid submissions before the deadline, and one team provided a valid submission after the deadline. All the submitted challenge results for each subset are presented in Tables 4-7. We noticed that if the pose estimation fails in one frame, it will cause an outlier with a large translation and rotation error. This makes the average translation and rotation error of several methods heavily affected by these extreme error values. Therefore, to rank the competing methods, we mainly focus on the ADD, Accuracy–ADD threshold Curve and the Avg Acc as defined in Section 3.2. IGTUM was awarded first place, and ICL was awarded second place. In the results, late submissions have been denoted using an asterisk.
Table 4: Evaluation on LND Test Without Occlusion subset. Best 2 methods are in bold.
Team ADD Avg Acc Translation Rotation
(10% diameter) ↑ (0-5 MM)↑ Error (mm)↓ Error (degree) ↓
@@
UOL* 0.0161 0.0640 8.1701 15.3407
PVNet** 0.1930 0.2866 46.7894 52.4488
    post-challenge submission; Avg Acc: Average Accuracy;
    baseline method.
Table 5: Evaluation on MBF Test Without Occlusion subset. Best 2 methods are in bold.
Team ADD Avg Acc Translation Rotation
(10% diameter) ↑ (0-5 MM)↑ Error (mm)↓ Error (degree) ↓
@@
UOL* 0.0239 0.0966 6.8555 10.4184
PVNet** 0.3589 0.4198 3.5265 25.8501
    post-challenge submission; Avg Acc: Average Accuracy;
    baseline method.
@@ 7.1 Instrument Pose Estimation Without Occlusion
The instrument pose estimation performance without occlusion was evaluated across two test subsets namely, LND Test Without Occlusion and MBF Test Without Occlusion using the ADD (10% diameter), Avg Acc (0-5 MM), Translation Error, and Rotation Error as shown in Table 4 and Table 5.
Fig. 9(a) and Fig. 9(b) illustrate the Accuracy curve with respect to different ADD thresholds for the examined subsets. It can be noticed that the accuracy of IGTUM has a significant gap from the other compared methods for the subset LND Test Without Occlusion while IGTUM, ICL, MVL_3S and PVNet have comparable accuracy for the MBF Test Without Occlusion subset.
@@
Figure 9: Accuracy–ADD threshold Curves for test data in Testsets Without Occlusion.
@@ 7.2 Instrument Tracking with Occlusion
Considering that pose estimation with occlusion is quite challenging, the performance of all methods under occlusion shows a significant drop as shown in Tables 6 and Tables 7. Across the LND Test With Occlusion and MBF Test With Occlusion subsets, IGTUM consistently outperformed the other methods with the highest ADD and Avg Acc scores and the lowest translation and rotation errors. This confirms its robustness and reliability under challenging conditions, such as occlusions. ICL, EUT and PVNet also showed promising performance in terms of ADD and Avg Acc scores, but with higher translation and rotation errors than IGTUM, indicating their limited effectiveness in challenging scenarios. TUDU and MVL_3S had lower ADD and Avg Acc scores across both subsets, coupled with moderate to high translation and rotation error rates, suggesting that these methods are less reliable under occlusion. The above performance trends are verified by the Accuracy-ADD threshold curves shown in Fig. 10(a) and Fig. 10(b).
Overall, for both cases with and without occlusion, IGTUM outperformed the other methods, consistently achieving the highest accuracy (ADD and Avg Acc) and the lowest errors (Translation and Rotation) across all four subsets, indicating its robustness and effectiveness. It achieved 36.06% ADD on average and 45.48% Avg Acc for all subsets. MVL_3S demonstrated good performance in the MBF Test Without Occlusion subset (Table 5), with competitive ADD and Avg Acc scores and overall lower translation and rotation errors. It achieved 17.19% ADD on average and 26.38% Avg Acc for all four subsets. TUDU exhibited moderate accuracy overall with 10.09% ADD on average and 20.15% Avg Acc for all subsets. UOL achieved 4.04% ADD on average and 14.63% Avg Acc for all four subsets. To be noticed, MVL_3S and TUDU achieved low translation errors but higher rotation errors. ICL, EUT and PVNet demonstrated strong performance in terms of ADD and Avg Acc scores but higher errors in translation and rotation error. EUT achieved 17.94% ADD on average and 24.06% Avg Acc while ICL obtained 22.92% ADD on average and 29.52% Avg Acc for all subsets.
The performance of the compared methods has also been evaluated on the secondary validation metrics, namely proj2d in Table 8 and mmd5 in Table 9 on all four subsets. IGTUM still outperformed the other methods with the best proj2d and mmd5 scores. ICL, EUT and PVNet demonstrated the second-best performance in the group since they followed a similar pipeline. UOL, TUDU and MVL_3S had lower proj2d and mmd5 scores across both subsets.
@@
Figure 10: Accuracy–ADD threshold Curves for test data in Testsets With Occlusion.
Table 6: Evaluation on LND Test With Occlusion. Best 2 methods are in bold.
Team ADD Avg Acc Translation Rotation
(10% diameter) ↑ (0-5 MM)↑ Error (mm)↓ Error (degree) ↓
@@
UOL* 0.0378 0.1894 8.3454 17.7254
PVNet** 0.2731 0.3901 28.0907 17.5543
    post-challenge submission; Avg Acc: Average Accuracy;
    baseline method.
Table 7: Evaluation on MBF Test With Occlusion Best 2 methods. are in bold.
Team ADD Avg Acc Translation Rotation
(10% diameter) ↑ (0-5 MM)↑ Error (mm)↓ Error (degree) ↓
@@
UOL* 0.0078 0.0346 15.9003 21.9342
PVNet** 0.1731 0.1826 44.6732 31.0920
    post-challenge submission; Avg Acc: Average Accuracy;
    baseline method.
Table 8: Evaluation on proj2d. Best 2 methods are in bold.
Team LND Test MBF Test LND Test MBF Test
w/o Occlusion w/o Occlusion w Occlusion w Occlusion
@@
UOL* 0.0080 0.0042 0.0143 0.0129
PVNet** 0.5308 0.8182 0.3487 0.5607
    post-challenge submission;
    baseline method.
Table 9: Evaluation on mmd5. Best 2 methods are in bold.
Team LND Test MBF Test LND Test MBF Test
w/o Occlusion w/o Occlusion w Occlusion w Occlusion
@@
UOL* 0.0134 0.0126 0.0718 0.0103
PVNet** 0.1903 0.0504 0.4928 0.1757
    post-challenge submission;
    baseline method.